# BulletGen: Improving 4D Reconstruction with Bullet-Time Generation

### v1

@@ 3.2 Generative augmentation
Time selection. One of the hyperparameters of our method is the number $n_{S}$ of sampled bullet time stamps. We sample them uniformly between the first and last time stamps, i.e. $t=1$ and $t=N$. In order to guarantee smooth motion, it is important to keep some distance between sampled time stamps since each bullet time generation may contain slight inconsistencies. If the distance between two sampled time stamps is low, this can generate effects similar to jittering. However, if the distance is sufficiently large, it gives enough time for smooth transitions between two bullet-time reconstructions with their own slight inconsistencies. For robust estimation, we start with the middle frame bullet-time stamp, e.g. $N/2$. Then, we proceed with the furthest time stamps. For instance, when $N=100$ and $n_{S}=5$, the sampled time stamps are $\{1,25,50,75,100\}$, but in the following order $\{50,1,100,25,75\}$. Also note that the bullet-time stamp selection can be progressive and proceed auto-regressively, and in our example, can continue with additional time stamps $\{12,37,62,87\}$ for $n_{S}=9$.
@@
View selection. The used generative model currently supports only static scenes and comes with two operating modes for generating left- and upward trajectories. By flipping the image horizontally, applying the leftward model, and flipping back, we can also generate rightward trajectories. Note that we do not apply the same principle to the upward direction, because the model only works well with upright imagery. Thus, we select between three modes for the view selection, i.e. left, right, up. Once bullet time is selected, we proceed with $n_{G}$ generations for the selected time. We tested three modes. First, for $n_{G}=3$, we select {up, left, right} in this order. For $n_{G}=5$, we select {up, left, right, left, right}, and for $n_{G}=7$, the used sequence is {up, left, up, right, up, left, right}. The view selection for generative model conditioning for each chosen direction is as follows. For leftward motion, we choose the left-most image (either the original one or the generated one), where the angle is measured from the middle of the original camera poses and viewing at the middle of the dynamic scene. Similarly, for rightward motion, we choose the right-most image. In all cases, only images at the currently selected bullet-time stamp are considered, either generated or the original one (there is only one original image for every time stamp). For upward motion, we choose the last generated image at this bullet time (or the original one if this is the first generation).
@@
Figure 4: Qualitative evaluation on several benchmark datasets. The input is a monocular video as shown in the training view column. Both benchmark datasets (Nvidia [87] and DyCheck iPhone [14]) have several additional testing cameras. We compare novel view synthesis to Shape-of-Motion (SoM) [73] and zoom-in to highlight the differences. Our method is able to provide more accurate and sharper reconstructions, both in static and dynamic parts of the scene.
Implementation details. All optimizations are performed using the ADAM optimizer [29]. The batch size for the generative views and the original video frames is set to 8 (thus, we have two concatenated batches of 8 in every iteration). The hyperparameters in Eq. (1) have been empirically determined and fixed to $\alpha_{1}=0.02$, $\alpha_{2}=0.1$, $\alpha_{3}=0.1$, $\alpha_{4}=0.5$ for all experiments. The threshold for generative views is set to $\gamma=0.4$. For the SoM loss function on the original video, we use their default weights. The total number of generated views per generation is fixed at $K=50$. The number of generations used is $n_{G}=7$, while the number of selected bullet times is $n_{S}=9$, unless specified otherwise. The average optimization time on a sequence from the iPhone dataset [14] with full resolution takes around 3 hours (including initial SoM optimization of 1.5 hours) on an Nvidia A100 80GB GPU. All main experiments required around 1 week of A100 GPU time, and preliminary non-included experiments required an additional month of GPU time. For evaluation on custom data, we use MegaSaM [38] to estimate camera poses. For evaluation on benchmark datasets, we used provided poses computed from COLMAP [63, 64] for fair comparison to other methods.
@@ 4 Experiments
We evaluate on the DyCheck iPhone dataset [14] and Nvidia dynamic dataset [87], which are commonly used datasets for novel view synthesis evaluation in dynamic scenes.
View synthesis metrics. We evaluate the standard PSNR, SSIM, and LPIPS scores between the ground truth and the rendered novel views. On the iPhone dataset, we measure those metrics on the covisible regions for fair comparison to other methods. As mentioned by [42], these scores are not robust enough on the current datasets due to camera misalignment, color differences between the training and test cameras, as well as the nature of the task. Thus, as proposed in [42], we also report CLIP-I [60] scores, which we see as more representative in this ill-posed task.
Tracking metrics. The iPhone dataset contains long-range 3D point tracking annotations. Thus, we also measure 3D end-point-error (EPE) at every timestep and the percentage of points that are within a certain radius of the ground truth 3D point (denoted by $\delta_{3D}^{\tau}$, where $\tau$ is either 5cm or 10cm). By projecting those 3D tracking annotations, we obtain 2D tracking annotations and report Average Jaccard (AJ) metric, average position accuracy ($<\delta_{avg}$) and Occlusion Accuracy (OA) metric.
Baselines. We compare to a range of methods that are specifically designed for dynamic scene reconstruction. Methods based on Neural Radiance Fields (NeRF) [51] and its variants are T-NeRF [35], HyperNeRF [56], and DynIBaR [39]. More recently, methods based on Gaussian Splatting were proposed, such as Deformable-3DGS [84], 4D-GS [76], Shape-of-Motion [73], and HiMoR [42]. We also compare to methods based on generative diffusion models, such as CoCoCo [96], StereoCrafter [93], ViewCrafter [88], CAT4D [77], and Vivid4D [22]. Many of these methods are concurrent with our work, are only pre-prints (not published), and do not provide code, e.g. [77, 22, 42]. Nevertheless, we strive to provide insightful comparisons by evaluating on the reported datasets chosen by these methods. For 2D and 3D tracking scores, we additionally compare to CoTracker [24] and TAPIR [11], which are lifted to 3D by DepthAnything [83].
@@
HyperNeRF [56] 0.182 28.4 45.8 10.1 19.3 52.0
DynIBaR [39] 0.252 11.4 24.6 5.4 8.7 37.7
Deformable-3D-GS [84] 0.151 33.4 55.3 14.0 20.9 63.9
CoTracker [24] + DA [83] 0.202 34.3 57.9 24.1 33.9 73.0
TAPIR [11] + DA [83] 0.114 38.1 63.2 27.8 41.5 67.4
Shape-of-Motion [73] 0.082 43.0 73.3 34.4 47.0 86.6
@@
T-NeRF [35] 15.60 0.55 0.55 0.86
HyperNeRF [56] 15.99 0.59 0.51 0.87
Deform.-3DGS [84] 11.92 0.49 0.66 0.79
SoM [73] 16.72 0.63 0.45 0.86
HiMoR [42] (no code) - - 0.46 0.89
CAT4D [77] (no code) 17.39 0.61 0.34 -
@@
Table 2: Evaluation on the iPhone dataset, novel view synthesis. The proposed method achieves state-of-the-art performance in terms of novel view synthesis.
@@
4D-GS [76] 14.01 0.39 0.59
CoCoCo [96] 14.99 0.47 0.53
StereoCrafter [93] 14.85 0.49 0.57
ViewCrafter [88] 14.94 0.49 0.58
SoM [73] 14.56 0.46 0.53
Vivid4D [22] 15.20 0.50 0.49
@@
Table 3: Evaluation on the iPhone dataset, subset chosen by Vivid4d [22]. Our method outperforms all other methods.
The iPhone dataset. The iPhone dataset [14] contains 12 sequences with a moving training camera captured with hand-held iPhone device, two static test cameras (for 5 sequences), and 3D tracking annotations. As shown in Table 1, our method outperforms all other state-of-the-art methods on all metrics in terms of 2D and 3D tracking scores. This is achieved by providing more constraints for scene geometry, which allows for better 3D triangulation during scene reconstruction. We also achieve state-of-the-art scores in terms of novel view synthesis (Table 3). For the sake of completeness, we add results reported in CAT4D [77] (unpublished work), which achieves slightly better PSNR and LPIPS scores, whereas our method obtains better SSIM results. However, we want to stress that the table remains incomplete in terms of CLIP-I score due to a lack of publicly accessible code and missing information about the training procedure and its overall training scale. To compare to the recently released Vivid4D [22] method (also without publicly available code), we choose their subset of the iPhone dataset that was used in their evaluation. Our proposed method again outperforms other methods (Table 3).
@@
Table 4: Evaluation on the Nvidia dataset. Our method improves the Shape-of-Motion (SoM) reconstruction on all metrics.
The Nvidia dataset. The Nvidia dataset contains 9 dynamic scenes captured by 12 synchronized static cameras. We use camera 1 as input, and cameras 2 and 3 as output for evaluation. This dataset contains testing views close to the training views, therefore the CLIP-I and LPIPS metrics do not show significant improvement, but in terms of PSNR, our improvement w.r.t. Shape-of-Motion is significant (see Table 4).
@@
Ablation study. The main hyperparameters of our method are the number of generations $n_{G}$ per selected bullet-time stamp, and the total number of those time stamps $n_{S}$. Table 5 shows that increasing $n_{G}$ per time stamp slightly improves most metrics, but is not significant given that the test cameras are not far from the training ones. A high value for $n_{G}$ is important for extreme novel view point synthesis. However, the number of selected bullet time stamps $n_{S}$ is vital even in the standard setting, and a higher number of time stamps improves all metrics.
@@
Figure 5: Temporal plane slices of extreme camera views. Visualizing the highlighted rows across time in $xt$ space shows that our method suffers from fewer temporal artifacts than Shape-of-Motion (SoM) caused by floating Gaussians and exploding geometry. Our rendered view is shown on the left.

### v2

@@ 3.2 Generative augmentation
Time selection. One of the hyperparameters of our method is the number $n_{S}$ of sampled bullet-time stamps. We sample them uniformly between the first and last time stamps, i.e. $1$ and $N$. For robust estimation, we start with the middle frame bullet-time stamp, e.g. $N/2$. Then, we proceed with the furthest time stamps. For example, when $N=100$ and $n_{S}=5$, the sampled time stamps are $\{1,25,50,75,100\}$, but in the order $\{50,1,100,25,75\}$. Also note that the bullet-time stamp selection can be progressive and proceed auto-regressively, and in our example, can continue with additional time stamps $\{12,37,62,87\}$ for $n_{S}=9$.
@@
View selection. The used generative model supports only static scenes and comes with two operating modes for generating left- and upward trajectories. By flipping the image horizontally, applying the leftward model, and flipping back, we can also generate rightward trajectories. Note that we do not apply the same principle to the upward direction, because the model works only well with upright imagery. Thus, we select between three modes for view selection, i.e. left, right, up. Once the bullet time is selected, we proceed with $n_{G}$ generations for the selected time. We tested three modes. First, for $n_{G}=3$, we select {up, left, right} in this order. For $n_{G}=5$, we select {up, left, right, left, right}, and for $n_{G}=7$, the used sequence is {up, left, up, right, up, left, right}. The view selection for the generative model conditioning for each chosen direction is as follows. For leftward motion, we choose the left-most image (either the original one or the generated one), where the angle is measured from the middle of the original camera poses and viewing at the middle of the dynamic scene. Similarly, for the rightward motion, we choose the right-most image. In all cases, only images at the currently selected bullet-time stamp are considered, either generated or the original (there is only one original image for every time stamp). For upward motion, we choose the last generated image at this bullet time (or the original one if this is the first generation).
@@
Figure 4: Qualitative evaluation on several benchmark datasets. The input is a monocular video as shown in the training view column. Both benchmark datasets (Nvidia and DyCheck iPhone) have several additional testing cameras. We compare novel view synthesis to Shape-of-Motion and zoom-in to highlight the differences. Our method is able to provide more accurate and sharper reconstructions, both in static and dynamic parts of the scene.
Implementation details. All optimizations are performed using the ADAM optimizer [29]. The batch size for the generative views and the original video frames is set to 8 (that is, we have two concatenated batches of 8 in every iteration). The hyperparameters in Eq. (1) have been empirically determined and fixed to $\alpha_{1}=0.02$, $\alpha_{2}=0.1$, $\alpha_{3}=0.1$, $\alpha_{4}=0.5$ for all experiments. The threshold for generative views is set to $\gamma=0.4$. For the SoM loss function on the original video, we use their default weights. The total number of generated views per generation is fixed at $K=50$. The number of generations used is $n_{G}=7$, while the number of selected bullet times is $n_{S}=9$, unless otherwise specified. The average optimization time on a sequence from the iPhone dataset with full resolution takes around 3 hours (including initial SoM optimization of 1.5 hours) on an Nvidia A100 80GB GPU. For evaluation on custom data, we use MegaSaM [38] to estimate camera poses. For evaluation on benchmark datasets, we used provided poses computed from COLMAP [64, 65] for a fair comparison to other methods.
@@ 4 Experiments
We evaluate on the DyCheck iPhone [14] and Nvidia dynamic [89] datasets, which are commonly used datasets for novel view synthesis evaluation in dynamic scenes.
View synthesis metrics. We evaluate the standard PSNR, SSIM, and LPIPS scores between ground truth and novel rendered views. On the iPhone dataset, we measure those metrics on the covisible regions for a fair comparison to other methods. As mentioned by Liang et al. [41], these scores are not robust enough in current datasets due to camera misalignment, color differences between training and test cameras, as well as the nature of the task. Thus, we also report CLIP-I [61] scores, which we see as more representative in this ill-posed task.
@@
HyperNeRF 0.182 28.4 45.8 10.1 19.3 52.0
DynIBaR 0.252 11.4 24.6 5.4 8.7 37.7
Deform.-3D-GS 0.151 33.4 55.3 14.0 20.9 63.9
CoTracker + DA 0.202 34.3 57.9 24.1 33.9 73.0
TAPIR + DA 0.114 38.1 63.2 27.8 41.5 67.4
Shape-of-Motion 0.082 43.0 73.3 34.4 47.0 86.6
@@
T-NeRF 15.60 0.55 0.55 0.86
HyperNeRF 15.99 0.59 0.51 0.87
Deform.-3DGS 11.92 0.49 0.66 0.79
Shape-of-Motion 16.72 0.63 0.45 0.86
HiMoR - - 0.46 0.89
CAT4D (no code) 17.39 0.61 0.34 -
@@
Table 2: Evaluation on the iPhone dataset. BulletGen achieves state-of-the-art performance in novel view synthesis.
@@
4D-GS 14.01 0.39 0.59
CoCoCo 14.99 0.47 0.53
StereoCrafter 14.85 0.49 0.57
ViewCrafter 14.94 0.49 0.58
Shape-of-Motion 14.56 0.46 0.53
Vivid4D (no code) 15.20 0.50 0.49
@@
Table 3: Evaluation on the iPhone dataset, subset chosen by Vivid4d. Our method outperforms all other methods.
Tracking metrics. The iPhone dataset contains long-range 3D point tracking annotations. Thus, we also measure 3D end-point error (EPE) at every timestep and the percentage of points that are within a certain radius of the ground truth 3D point (denoted by $\delta_{3D}^{\tau}$, where $\tau$ is either 5cm or 10cm). By projecting those 3D tracking annotations, we obtain 2D tracking annotations and report Average Jaccard (AJ) metric, average position accuracy ($\delta_{avg}$) and Occlusion Accuracy (OA) metric.
Baselines. We compare to a range of methods that are specifically designed for dynamic scene reconstruction. Methods based on Neural Radiance Fields (NeRF) [51] and its variants are T-NeRF [35], HyperNeRF [57], and DynIBaR [39]. More recently, methods based on Gaussian Splatting were proposed, such as Deformable-3DGS [87], 4D-GS [77], Shape-of-Motion [74], and HiMoR [41]. We also compare to methods based on generative diffusion models, such as CoCoCo [98], StereoCrafter [95], ViewCrafter [90], CAT4D [79], and Vivid4D [22]. Some methods do not provide code, e.g. CAT4D and Vivid4D. Nevertheless, we strive to provide insightful comparisons by evaluating on the reported datasets chosen by these methods. For 2D and 3D tracking scores, we additionally compare to CoTracker [24] and TAPIR [11], which are lifted to 3D by DepthAnything [85].
Figure 5: Temporal plane slices of extreme camera views. Visualizing the highlighted rows across time in xtxt space shows that our method suffers from fewer temporal artifacts than Shape-of-Motion (SoM) caused by floating Gaussians and exploding geometry. Our rendered view is shown on the left.
The iPhone dataset. The iPhone dataset contains 12 sequences with a moving training camera captured with a hand-held iPhone device, two static test cameras (for 5 sequences), and 3D tracking annotations. As shown in Table 1, our method outperforms all other state-of-the-art methods on all metrics in terms of 2D and 3D tracking scores. This is achieved by providing more constraints for scene geometry, which allows for better 3D triangulation during scene reconstruction. We also achieve state-of-the-art scores in terms of novel view synthesis (Table 2). We also add results reported in CAT4D (no code), which achieves slightly better PSNR and LPIPS scores, whereas our method obtains better SSIM results. However, we want to stress that the table remains incomplete in terms of CLIP-I score due to a lack of publicly accessible code and missing information about the training procedure and its overall training scale. To compare to Vivid4D, we choose their challenging subset of the iPhone dataset. Our method again outperforms other methods (Table 3). This highlights that our method is more effective in those challenging cases (as compared to the full dataset).
@@
Table 4: Evaluation on the Nvidia dataset. Our method improves the Shape-of-Motion reconstruction on all metrics.
@@
The Nvidia dataset. The Nvidia dataset contains 9 dynamic scenes captured by 12 synchronized static cameras. We use camera 1 as input, and cameras 2 and 3 as output for evaluation. This dataset contains testing views close to the training views, therefore the CLIP-I and LPIPS metrics do not show a significant improvement, but in terms of PSNR, our improvement w.r.t. SoM is significant (Table 4).
Ablation study. The main hyperparameters of our method are the number of generations $n_{G}$ per selected bullet-time stamp, and the total number of those time stamps $n_{S}$. Table 5 shows that increasing $n_{G}$ per time stamp slightly improves most metrics, but the improvement is not significant given that the test cameras are not far from the training ones. A high value for $n_{G}$ is important for extreme novel view point synthesis. However, the number of selected bullet time stamps $n_{S}$ is vital even in the standard setting, and a higher number of time stamps improves all metrics.